OpenStudy (anonymous):

Mrs. Manning's current savings is 156% of what she had in her account last month. If she has $1170 in her account this month, what was her last month's savings?

OpenStudy (triciaal):

last month = x 156% x = 1170 x = 156% x /156% = 1170/ 156% 1170/1.56 = 750

OpenStudy (triciaal):

1170 - 750 = 420 new balance - old balance = amount saved